the sum of all natural numbers less than 200 that are divisible by 3 but not by 5 is

Grade:11

1 Answers

atulya yadav
6 years ago
Sum of numbers less than 200 and divisible by 3 is3+6+9+...........198=66(3+198)/2=6633Now in this sum there are also numbers like 15 ,30,45 which is divisible by 5 so subtract it 15+30+45+.........195=13 (15+195)/2=1365 so answer is 6633-1365=5268